In this page, you will find all the manuals needed to prepare and analyze your observations with CRIRES. Please be sure to download that corresponding to the period in which the observations will take place.
Note that CRIRES was removed from UT1 in July 2014 for an upgrade (new detectors, cross-disperser, calibration sources). It was not be offered from P94 onwards. It is expected that CRIRES plus will be back on Paranal during mid-2018. For more information see the CRIRES plus webpages.
All the manuals are in Adobe's Acrobat format.
- CRIRES User's Manual
All the information about the instrument itself, its modes and their characteristics:
- for Period 93, Phase 2 (v93.2)
- for Period 92, Phase 2 (v92.2)
- for Period 91, Phase 2 (v91.2)
- for Period 90, Phase 1 (v90.2)
- for Period 89, Phase 2 (v89.2)
- for Period 88, Phase 2 (v88.4)
- for Period 87, Phase 2 (v87.3)
- for Period 86, Phase 2 (v86.2)
- for Period 85, Phase 2 (v85.2)
- for Period 84, Phase 2 (v84.2)
- for Period 83, Phase 2 (v83.2)
- Science Data Product Forum
A forum http://www.eso.org/sci/data-processing/forum.html is available to exchange ideas, methods, software
and data to assist with the production of science data products from ESO telescopes.
- CRIRES data reduction frequently asked questions
A list of Frequently Asked Questions http://www.eso.org/sci/data-processing/faq.html
on data reduction is also available.
- CRIRES descriptions
- Description of CRIRES Käufl et al. 2004, SPIE 5492, 1218
- Description of the MACAO used for CRIRES Paufiqué et al. 2004, SPIE 5490, 216
- The CRIRES InSb megapixel focal plane array detector mosaic Dorn et al. 2006, SPIE 5499, 510
- Other useful documents
- Interpixel capacitance in large format CMOS hybrid arrays Finger et al. 2006, SPIE 6276, 62760F
- History of CRIRES
Here we report changes that can affect the quality of thedata since start of operations
- Aug 28 2013: The metrology system has still not been implemented although it is hoped that this will be finally offered in P94. It is likely that in P94 the main wavelength calibration system will change from ThAr to UNe as this produces 2 to 5 times more lines in most wavelength regions.
- May 05 2013: The ThAr lamp failed and was replaced by a new model that has fewer emission lines, hence decreasing the quality of the wavelength calibration.
- Jun 22 2011: It is likely that with the change to fixed slits of 0.2 and 0.4 arcseconds and the introduction of metrology in P88 that the repeatability and accuracy of flatfields and wavelength calibrations will improve. To reduce the calibration load then the validity of calibrations may be increased from 1 day starting in P89.
- Jun 21 2011: The CRIRES ETC has been changed to (1) Decrease the predicted S/N ratio by a factor 1.3 for wavelengths beneath 1.1 microns (2) Only allow seeing of less than 1.4 arcseconds for AO OBs.
- Apr 14 2011. Dark levels for 300-s DIT darks now back to normal: Jun 22 2011: CRIRES user manual was released on June 20th 2011 for P88 phase 2.
- Mar 27 2011. Detector 4 is showing some features in the 300-s darks on the right hand side of the detector, and the readout noise has slightly increased for detectors 3 and 4. Under investigation.See here for an example of detector 4 300-s dark taken on March 25th 2011.
- Mar 20 2011. UTC Time: 17:11:59 there was an earthquake thatshifted the prism, hence moving the wavlength zero-point for some ofthe calibrations. The situation was recovered on March 21st and noscience data were affected.
- Mar 08 to 16th 2011. An intervention on CRIRES was performedon one of the closed cycle coolers. CRIRES back in operations on March16th.
- Feb 18, 2011. An intervention is planned in early March 2011 for maintenance of the closed cycle coolers.
- Nov 27, 2010. CRIRES back in operations.
- Nov 26, 2010. The warm part of CRIRES had to be dismounted onNov 13 due to M1 and M3 recoating. The instrument should be back soon.
- Sept 07, 2010. CRIRES is back in operations. The X-positionof the slit on the slit viewer detector has changed from 532 to 568pixels that means that the field of view to the right of the slit hasdecreased by about 2 arcseconds. The allowed deltaX range is now -17.5"and 16.5". The slit is parallel to 0.02 arcseconds for a 10 arcsecondthrow and 0.04 arcseconds for a 20 arcsecond throw. User manual updatedon 16-Sept-2010 to reflect the changes above.
- Aug 22, 2010. Another intervention in progress.
- Aug 21, 2010. Problems with service mode CRIRES data are noted in the CRIRES SM pages.
- Jul 16, 2010. Another intervention was performed that left a problem with the alignment of the entrance and intermediate slit. This means that only data from detector 2 and 3 will be useable for large nod-throws/short wavelengths. The slit is now parallel to better than 10 percent over a throw of 20 arcseconds. The flatfields no longer show gradients in the cross dispersion.
- 2010-05-01 until 2010-07-10. There were more problems with CRIRES cooling that meant that the instrument was not available for the majority of May as a new cooling system needed to be installed. After this intervention there were found to be the following problems that will affect data quality: (1) The slit shows 50 percent variation in its width from the top to bottom, and 20 percent variation if a nod-throw of 10 arcseconds is used. (2) In the cross-dispersion direction the flatfields show a gradient of nearly a factor 3. An intervention is planned in early July to try and fix these issues.
- 2010-02-22: Intervention on closed cycle coolers in order to improve the closed cycle cooling.
- 2009-12-25: Metrology fibres focussed - when implemented in time for P86 these should improve the wavelength stability of the instrument for standard settings only. Due to the intervention the position of lines on the grating shifted by between 10 and 30 pixels in the wavelength direction and 5 pixels in the cross-dispersion direction. Intervention to try and find the leak in the closed cycle coolers that was causing contamination of the detectors that meant bi-monthy partial warmups of the instrument were necessary.
- 2009-05-31/06-01: SW: (a) Because of the change of the way the intermediate slit works (one jaw blocked), the spectrograph configuration table had to be modified. (b) Since the internal functions are now well reproducible, the algorithm used since the start of operations to insure that the spectrum is well centered in the intermediate slit has now been removed. Overheads for a change of wavelength setting are therefore reduced significantly. In December 2009, increased reproducibility will be insured by the use of the metrology fibers. (c) Differential tracking is now fully implemented.
- 2009-05 : Intervention: (a) the intermediate slit mechanism was replaced by a much more reliable one. One slit jaw is fixed. The uncertainty on the slit width is now a few microns. An initial test after the instrument was first cooled revealed an electrical short that forced a re-opening. Initial measurements show that the reproducibility of the illumination in the non-contaminated area is much below 1%; (b) the science detector mosaic was moved again, this time in the correct direction. Its center is closer to the optical axis. The effective slit length has increased by 30%: it is now slightly larger than 40 arcsec. The change also brings the optical axis closer to the center of the slit viewer detector therefore reducing the load on the dynamics of the MACAO tip-tilt mount (only about 20% compared to nearly 60%). (c) Metrology fibers were installed both at the level of the entrance slit and at the level of the intermediate slit. Once the project is completed, they would ensure reproducibility of the wavelength setting at the sub-pixel level. As expected, a fine adjustment is needed for the entrance slit width fibers (focus, position) which will be carried out in a new intervention planned for December 2009.A side-effect of the installation of the metrology fibers is that ThAr frames not only show lines from the wavelength calibration fibers but also from the metrology fibers: UM shows an example; (d) ne of the CCCs is ran at somewhat higher speed to increase the cooling power. Also, the set point for the temperature for various parts of the instrument has been incrased slightly. These 2 changes should allow a better thermal control; (e) The J band filter is still not useable.
- 2009-02-03/07: The CRIRES SW v2.0 installation is now completed. The SW is much more stable with little down time. New functionalities include full implementation of differential effects between the different wavelength involved: differential refraction, chromatic effect of entrance window. These effects are re-evaluated regularly, including after change of wavelength settings, change ofairmass, change of SV filter. In addition, the 1st phase of the implementation of differential tracking was carried out.
- 2009-01 : Intervention: (a) the cryo-coolers (CCC) were changed by models that do not introduce vibrations impeding observations of the VLTI. As a consequence, the CRIRES CCC do not have to be stopped anymore during VLTI UT runs. (b) the science detector was moved along the spatial direction; the intention was to increase the slit length but the unfortunately, the illuminated area is closer to the regions where the detector glows are located; (c) re-cabling allowed to strongly decrease the noise affecting all the encoders; they now behave within or close to specifications, although some excess noise is still present; (d) the J band filter was changed to bring images to focus; however, the problem persists; (e) a patch was introduced in SW to deactivate the motors (grating, prism) at the end of the motor movement, therefore reducing the jitter. caused by a small target radius and the residual jitter of the resolver.
- 2008-11-18: The MACAO safety shutter failed on Nov 18, 2008. Its replacement unfortunately led to a mis-alignment of the fibrebundle. After a week of efforts, its re-alignment wascompleted on Nov 27.
- 2008-10-29: After, a medium-risk earthquake, the grating encoder became noisy. Subsequent investigations indicate that the problem originates inside the vessel: therefore, it cannot be solved before the January 2009intervention. The spectral resolution is not affected. However, the wavelength setting is affected with a precision reduced to +/- 5 pixels.
- 2008-10-26: The 'noise' affecting the PRISM encoder has been reduced to asatisfactory value. Complete elimination shall wait for the next intervention in January 2009.
- 2008-08-13: Wavelength calibration using the ThAr fibers are now obtained with 4 fibers on the field (a 5th one is just on the edge of the field). Centering the fibers is carried out at each execution of the template as the light path through the MACAO is likely temperature dependent. Also, the DIT and baffle location are adjusted to provide optimal illumination of the lines while avoiding to have any pixel at more than 12,000 ADUs to avoid possible remanence effects.
- 2008-06-25: A new intervention took place this month: it successfully solved the problem of the vignetting of the entrance slit; it also solved a problem with the intermediate slit encoder. The instrument is in a cool-down phase (although interrupted by the VLTI run that requires the CCC to be off).
- 2008-05-25: During the warm-up/cool-down cycle associated with the VLTI run (required as CRIRES CCCs are causing too many vibrations for VLTI), the ThAr fiber connector has been temporarily replaced by a connector providing only 1 fiber on the unvignetted area, to allow for a small modification. Following this warm-up/cool-down cycle, the entrance slit vignetting has increased. Unvignetted area is now 8.1 arcsec. In the mean time, the cause of the vignetting has been identified as a detaching Sheldal tape when the instrument is below 80K and before it reached operational temperature. Intervention to solve the problem will take place on June 14. Restart of operations foreseen by June 28-29, after the June VLTI run. In the mean time, CRIRES operations are limited to OBs that cannot be observed after the June intervention. On a separate matter, the CRIRES_spec_cal_Wave, when used with the ThAr lamp, has been modified to make sure that the fiber is well centered in the slit: possible temperature effects on the MACAO optics lead to uncertain centering.
- 2008-04 : Hardware intervention: (i) carriage: replaced by a faster translation table; gas-cells now easily accessible, incl. CO gas-cell now available; AO fiber installed; ThAr lamp on the carriage removed, replaced by a series of 7 fibers behind a variable neutral density filter; only 2 pinholes kept; (ii) pre-optics reasonably shielded against dome-light; (iii) science detector now without vignettined; (iv) thermal control of the spectrograph considerably improved. Bad news: entrance slit shows vignetting with initially unidentified origin. At the start of operations, unvignetted slit-length is 12 arcsec. Because of the vignetting, only 2 ThAr fibers can be seen in the field-of-view.
- 2007-12 : Hardware intervention: SV detector and entrance slit jaws have been replaced. Cryostat temperature set to 70 K (instead of 65 K), requiring a small change in the wavelength configuration table. Entrance slit is now set in open loop due to noisier encoder.
- 2007-08-06: Correct a bug that set columns 512 and 513 of detector 1 and 4 to 0 in the INT frames and in the FITS files.
- 2007-07-03: Algorithm added to all acquisition (AO, NoAO), observation (obs/cal_AutoNodOnSlit),and calibration templates. This algorithm adjusts the piezo andpossibly the prism orientation so that the spectrum is well centered in the intermediate slitand on the detector mosaic. This is a first step to work around theproblem of the intermediate slit lack of reproducibility, until this function is replaced.
- 2007-06-07: For FF and ThAr, minimum DIT is 2s.